# : Adaptive Image Focus Optimization for Black-box MLLM

### v1

@@ 5 Experiments
In this section, we evaluate the performance of through a series of experiments designed to test its ability to improve token efficiency and preserve visual fidelity across different black-box MLLM. Specifically, we aim to answer the following questions: (i) Accuuracy: Does improve accuracy across different black-box MLLMs on image-related tasks? (ii) Efficiency: How does perform compared to baseline methods in terms of both accuracy and token efficiency? (iii) Component Contribution: What is the impact of key components in , such as multi-scale vision emphasize and the budget-aware prompt strategy?
5.1 Setup
1) VstarΒ [28]: A benchmark dataset focused on image classification, used to evaluate fine-grained visual recognition capabilities in object detection and classification tasks.
2) CVBenchΒ [24]: Contains 2 sub-category, and , respectively, representing two-dimensional and three-dimensional visual image, respectively, to evaluate the performance of the model when processing images of different dimensions, especially the understanding ability in complex scenes.
3) RealworldQAΒ [29]: Used to test the multimodal question answering performance of the model in real-world scenarios, involving cross-language and cross-image information processing.
4) MMVPΒ [25]: A validation set for multimodal visual processing, designed to evaluate the comprehensive understanding of models for complex visual scenes.
5) ScienceQAΒ [18]: A multimodal scientific question-answering dataset featuring multiple-choice questions across a diverse range of science topics.
6) MMMUΒ [33]: The validation part of a new benchmark, which designed to evaluate the performance of multimodal models on multidisciplinary tasks that require university-level subject knowledge and deliberate reasoning.
7) HRΒ [27]: A high-resolution multimodal benchmark consisting of 4K and 8K images and corresponding questions.
@@ 5.2 Main results
TableΒ 2 compares the performance of against baseline methods across various datasets using GPT-4o. The results show that , particularly in its Patches and Adaptive versions, consistently outperformed baseline approaches in terms of accuracy while maintaining lower token usage than the Raw method. consistently outperformed baseline methods, showing accuracy improvements up to 26% across multiple tasks. For example, in the Vstar dataset, -Patches achieved an accuracy of 0.717, compared to 0.565 using the Raw baseline, marking a 26.9% improvement. In RealworldQA, which demands complex multimodal reasoning, -Adaptive achieved 0.758 accuracy, outperforming the 0.676 accuracy of the Raw method by 12.1%. These results highlight that effectively preserves fine-grained visual details, enabling improved object recognition and image understanding across real-world tasks, where precise detail retention is crucial.
We further evaluated on Claude-3.5-Sonnet and Gemini-1.5Pro to assess its generalizability across different black-box MLLMs. TableΒ 4 shows that demonstrated robust performance across different black-box MLLMs, for example, achieved an accuracy of 0.704 on Vstar, compared to 0.531 with the Raw baseline, marking a 32.6% improvement. Similarly, in Claude-3.5-Sonnet, outperformed the baseline by 34.5% on RealworldQA, improving from 0.610 to 0.741. These results suggest that can consistently enhance performance across different architectures, making it a versatile tool for various MLLM-based applications.
A key contribution of is its ability to reduce token consumption while maintaining or improving accuracy. As shown in TableΒ 3, consistently delivers both token efficiency and performance improvements across various benchmark datasets. For instance, on the TerraIncognita dataset, guided by ManyICLΒ [10], achieves 0.83 accuracy using 315 tokens, compared to the Raw baselineβs 0.78 accuracy with 963 tokensβa 67% reduction in token usage while improving performance by 6.4%. Additionally, reduces latency, making it practical for real-time applications. In the TerraIncognita zero-shot setting, lowered latency from 4.8s to 3.1s, a 35.4% reduction without sacrificing accuracy. This makes highly suitable for tasks like autonomous driving, and real-time visual analytics, where both token efficiency and reduced latency are critical.
@@ 5.3 Findings
Here we analyze why the -Patches version underperforms compared to -Global and even the Local version on certain datasets. For example, on the dataset, the accuracy of the Patches version is 0.025 lower than the Global version and 0.04 lower than the Local version. Similarly, on the MMVP dataset, the Patches version falls short by 0.005 compared to the Global version and by 0.003 compared to the Adaptive version. Given that these results are averaged over multiple measurements, and accounting for model fluctuations, we hypothesize that this performance drop occurs because the Patches version treats each RoI as an independent image and provides them separately to the MLLM. When there are too many RoIs, the model may fail to capture or integrate some of them, leading to a drop in accuracy.
@@ 5.4 Ablation study
To further understand the impact of the individual components within , we conducted an ablation study focusing on two key variants: 1) with multi-scale emphasize: Compared with the commonly used multi-resolution and directly use, multi-scale visual emphasis is used to identify and emphasize RoI in the image. 2) with different models.
To further investigate the contributions of βs components, we conducted an ablation study (TableΒ 5). The results demonstrate that the combination of multi-scale visual emphasis and Patches prompt strategies delivers the best performance across all most datasets. Comparing different vision emphasis models, such as EVF-SAM and Ground Dino, further highlights the effectiveness of . Despite differences in model capabilities, both show accuracy improvements across datasets. Additionally, when comparing different emphasis methodsβDefault, Multi-Resolution, and Multi-Scaleβthe Multi-Scale method consistently outperformed Multi-Resolution. We hypothesize that, while Multi-Scale crops images and may split objects, its pyramid-shaped multi-recall strategy compensates for this by enhancing recall. In contrast, although Multi-Resolution maintains object integrity, adjusting resolution disrupts the modelβs performance, likely because most models are trained on fixed-size inputs, and changing the resolution weakens their inherent capabilities.
@@ 6 Conclusion
In this paper, we introduced , a novel visual prompting mechanism designed to overcome the limitations of black-box MLLMs in processing images while adhering to token constraints. Our approach effectively balances the need to capture essential visual details without exceeding token budgets, a challenge commonly encountered in existing models like GPT-4o and Gemini Pro.
Through a comprehensive evaluation across datasets such as Vstar and RealWorldQA, demonstrated significant improvements, particularly in fine-grained visual tasks. Our results show that -Patches achieved a 26.9% accuracy gain over baseline methods in Vstar, and -Adaptive provided a 12.1% improvement in RealWorldQA. These gains were achieved while drastically reducing token usage, with delivering 6.4% higher accuracy in the TerraIncognita using 67% fewer tokens.
Although this work primarily focuses on improving the efficiency of visual processing in black-box MLLMs, another potential issue that arises in real-world applications is communication cost. For example, transferring large images from edge devices (e.g., wearable cameras or glasses) to cloud servers can be expensive in terms of bandwidth, latency, and energy consumption. While is designed to reduce token usage, its application in minimizing data transmission costs is an area that could be explored in future work.
As part of future research, we plan to investigate how could be adapted for edge ML applications, enabling local processing on devices such as wearable cameras. This would allow for more efficient handling of visual inputs at the edge, reducing the need for extensive data transfers to the cloud. We aim to measure latency, power consumption, and the overall impact on system performance to assess the feasibility of applying in these scenarios.
In summary, offers a practical solution for enhancing visual processing in constrained MLLMs and opens up new directions for future exploration.

### v2

@@ 4.2 Spatial-preserving Orchestration Schema
In all configurations, Zoomer composes the selected ROIs and optional global view into a single image canvas, which is then passed to the black-box MLLM. In particular, in Zoomer-Adaptive and Zoomer-Global, ROIs are placed according to their original relative coordinates using the spatial-preserving layout, and, when enabled, a down-scaled global overview is appended as an additional strip on the same canvas. Thus, the MLLM always receives exactly one reconstructed image rather than multiple separate images.
@@ 4.3 Budget-aware Prompting Strategy:
Adaptive global view decision.
Let EQΒ 1 denote the fraction of the original image covered by the selected ROIs. Zoomer-Adaptive decides whether to append a global overview based on : if , we append a down-scaled global view to preserve overall scene context; if , we omit the global view to save visual tokens, since the ROIs already cover a sufficiently large portion of the scene. The threshold is chosen empirically on the VSTAR validation set to balance accuracy and token efficiency.
@@ 5 Experiments
We evaluate the performance of on a diverse set of visual reasoning tasks using multiple black-box MLLMs. Our experiments are designed to assess both quantitative gains in accuracy and qualitative improvements in visual fidelity under constrained token budgets.
Specifically, our evaluation is guided by the following research questions: RQ1: Accuracy. Does improve task accuracy across different black-box MLLMs on image-grounded reasoning tasks, compared to standard prompting baselines? RQ2: Token Efficiency. How does balance the trade-off between accuracy and token usage, and how does it perform relative to methods that either over-utilize or under-utilize the token budget? RQ3: Component Effectiveness. What is the contribution of each major component of βspecifically, the multi-scale visual emphasis module and the budget-aware prompting strategyβto the overall performance?
Backbone MLLMs.
We evaluate Zoomer on both commercial APIs (e.g., GPT-4o, Gemini, Claude) and open-source MLLMs. For the latter, we use Qwen2-VL-7B and InternVL2.5-8B as representative models in this work, primarily due to time and resource constraints. As shown in TableΒ 6, applying Zoomer on top of both open-source backbones consistently improves performance across four benchmarks, indicating that the proposed adaptive visual token allocation is not tied to proprietary APIs. Extending the evaluation to additional high-resolution open-source models such as Qwen3-VL, Qwen2.5-VL, LLaVA-UHD, and MG-LLaVA is an important direction for future work.
5.1 Setup
1) VstarΒ (vstar): A benchmark dataset focused on image classification, used to evaluate fine-grained visual recognition capabilities in object detection and classification tasks.
2) CVBenchΒ (tong2024cambrian1): Contains 2 sub-category, and , respectively, representing two-dimensional and three-dimensional visual image, respectively, to evaluate the performance of the model when processing images of different dimensions, especially the understanding ability in complex scenes.
3) RealworldQAΒ (xAI2024): Used to test the multimodal question answering performance of the model in real-world scenarios, involving cross-language and cross-image information processing.
4) MMVPΒ (mmvp): A validation set for multimodal visual processing, designed to evaluate the comprehensive understanding of models for complex visual scenes.
5) ScienceQAΒ (sqa): A multimodal scientific question-answering dataset featuring multiple-choice questions across a diverse range of science topics.
6) MMMUΒ yue2023mmmu: The validation part of a new benchmark, which designed to evaluate the performance of multimodal models on multidisciplinary tasks that require university-level subject knowledge and deliberate reasoning.
7) HRΒ wang2024divideconquercombinetrainingfree: A high-resolution multimodal benchmark consisting of 4K and 8K images and corresponding questions.
@@ 5.1 Setup
Token efficiency metrics.
Different providers adopt different image tokenization schemes. Whenever an API exposes image-token usage, we directly report the providerβs visual token count. When this information is not exposed, we instead count the number of patches that are actually transmitted after reconstruction as a proxy for visual tokens.
To ensure fair cross-model comparison, we always measure token efficiency as a relative reduction with respect to each modelβs own Raw baseline. For a model and method , we compute
and pair this with the corresponding accuracy change. This normalization prevents differences in provider-specific tokenization rules from biasing our conclusions about token savings.
@@ 5.2 Main results
RQ1: Accuracy Under Token Constraints
We first evaluate the accuracy of across multiple visual reasoning benchmarks using GPT-4o. TableΒ 2 compares its performance with baseline prompting strategies, including Raw (unaltered image input), Crop, and heuristic patching. Across all datasets, demonstrates consistent and significant gains.
Notably, -Patches achieves 71.7% accuracy on the Vstar dataset, outperforming the Raw baseline by 26.9% (from 0.565 to 0.717). On the more challenging RealWorldQA dataset, which requires complex visual-linguistic reasoning, -Adaptive yields 0.758 accuracyβ12.1% higher than the Raw input (0.676). These results indicate that enhances the modelβs ability to ground object references and interpret fine-grained visual cues by preserving high-fidelity regions while respecting token limits.
RQ2: Cross-Model Generalizability
To assess generalizability, we apply to other closed-source MLLMs, including Claude-3.5-Sonnet and Gemini-1.5Pro. As shown in TableΒ 4, the performance trends hold across architectures. For instance, on Vstar, improves accuracy from 0.531 to 0.704 on Gemini-1.5Proβa 32.6% gain. On RealWorldQA, Claude-3.5-Sonnet achieves a 34.5% improvement using (from 0.610 to 0.741). These results suggest that is model-agnostic and portable across different commercial MLLMs without requiring model access or retraining.
As suggested by reviewers, we also evaluate Zoomer on open-source MLLMs. In particular, we use Qwen2-VL-7B and InternVL2.5-8B as representative backbones. As shown in TableΒ 6, applying Zoomer on top of both models yields consistent gains across all four benchmarks (e.g., +0.034β0.043 on RealWorldQA and +0.064β0.082 on VSTAR), confirming that our adaptive image focus mechanism is not tied to proprietary APIs.
| Qwen2-VL-7B | InternVL2.5-8B | Qwen2-VL-7B + Zoomer | InternVL2.5-8B + Zoomer | |
| CVBench-2D | 0.648 | 0.656 | 0.683 (+0.035) | 0.697 (+0.041) |
| CVBench-3D | 0.632 | 0.629 | 0.675 (+0.043) | 0.669 (+0.040) |
| RealWorldQA | 0.698 | 0.698 | 0.732 (+0.034) | 0.741 (+0.043) |
| VSTAR | 0.584 | 0.573 | 0.648 (+0.064) | 0.655 (+0.082) |
RQ3: Token Efficiency and Latency
One of the central goals of is to maximize information utility under token constraints. TableΒ 3 reports token usage and performance on the TerraIncognita dataset under the ManyICLΒ (manyICL) setting. achieves 0.83 accuracy with 315 tokens, while the Raw method reaches only 0.78 with 963 tokensβyielding a 6.4% accuracy improvement with 67% fewer tokens. This substantial reduction in visual token cost translates directly into lower API usage and faster inference.
Moreover, in real-time settings such as autonomous navigation, latency becomes critical. In a zero-shot TerraIncognita setting, reduces response time from 4.8 seconds to 3.1 seconds (a 35.4% reduction) without sacrificing accuracy. These results underscore the practical benefits of for latency-sensitive and resource-constrained applications such as real-time surveillance, embodied agents, and mobile vision systems.
We further compare Zoomer against a stronger high-resolution baseline Vβ, which already uses high-resolution visual inputs without our prompt-aware cropping and composition. As shown in TableΒ 7, Zoomer still provides clear gains, especially on CVBench-3D and RealWorldQA. This indicates that adaptive image focus and spatial-preserving orchestration remain beneficial even when the baseline already operates in a high-resolution regime, and strengthens our connection to high-resolution prompting methods such as LLaVA-UHD and MG-LLaVA.
| Method | VSTAR | CVBench-2D | CVBench-3D | RealWorldQA |
| Zoomer | 0.717(+0.013) | 0.746(+0.045) | 0.883(+0.090) | 0.758(+0.060) |
| Vβ | 0.704 | 0.701 | 0.793 | 0.698 |
@@ 5.3 Findings
While -Patches generally performs well under generous token budgets, we observe that in certain datasets it underperforms relative to other variants. For instance, on the dataset, -Patches yields an accuracy that is 2.5 percentage points lower than -Global and 4.0 points lower than -Local. Similarly, on the MMVP benchmark, the Patches variant trails -Global by 0.5 percentage points and -Adaptive by 0.3 points. These differences persist across multiple trials and are consistent beyond minor fluctuations attributable to model non-determinism.
We hypothesize that this degradation stems from the way -Patches encodes visual inputs. Specifically, each RoI is treated as an independent image patch and passed separately to the black-box MLLM. This approach, while maximizing local detail preservation, inherently disrupts the global spatial context and inhibits inter-region integration. When the number of RoIs increases, the model is forced to reason over fragmented visual inputs without spatial continuity, which may lead to failures in tasks requiring holistic scene understanding or relational reasoning.
Failure case: missed small object in ROI extraction.
Despite the overall gains, Zoomer may fail when the ROI extractor misses a small but question-critical object. FigureΒ 6 shows an example where the question asks: βIs the dog on the left or right side of the golden tower?β The dog is a tiny object in the scene. If the detector fails to propose an ROI covering the dog, Zoomer-Local will not include the dog in the reconstructed input, causing the MLLM to answer incorrectly. Zoomer-Adaptive may partially mitigate this issue by appending a global view, but it still cannot recover the missing evidence when the dog becomes indistinguishable in the down-scaled overview. This highlights a key limitation of detector-guided visual focus and motivates future detector-free visual focusing strategies.
These findings suggest that preserving global structureβeven at the cost of slightly reduced local detailβis beneficial in tasks that involve complex spatial or semantic dependencies. They also underscore the need for future designs to better balance granularity with contextual coherence in visual prompting pipelines.
@@ 5.4 Ablation Study
To assess the contribution of individual components in , we conducted a comprehensive ablation study along two primary dimensions: (1) the role of multi-scale visual emphasis strategies, and (2) the impact of different RoI localization models. The results are summarized in TableΒ 5.
Patch selection under a fixed budget.
As a stronger baseline than Raw and Resize, we compare Zoomer-Patches with Random and Uniform patch selection under the same patch budget (TableΒ 8). Since all three strategies transmit the same number of patches, the improvements of Zoomer-Patches can only come from semantically guided region selection and spatial-preserving composition, rather than from simply increasing visual resolution.
| Method | CVBench-2D | CVBench-3D | VSTAR |
| Zoomer-Patches | 0.746 | 0.858 | 0.717 |
| Random-Patches | 0.571(-0.175) | 0.613(-0.245) | 0.597(-0.120) |
| Uniform-Patches | 0.637(-0.109) | 0.731(-0.127) | 0.664(-0.047) |
Visual Emphasis Strategies.
We compare three emphasis methods: (i) a baseline strategy using the full image without emphasis ("Default"), (ii) a multi-resolution method that applies image resizing at various scales without cropping, and (iii) our proposed multi-scale strategy, which extracts overlapping RoI patches across hierarchical granularities.
Across all tested datasets, the multi-scale approach consistently outperforms the other two. For instance, on the Vstar and CVBench datasets, the multi-scale strategy yields relative improvements of 8.2% and 6.4% over the multi-resolution method, respectively. This suggests that the hierarchical region recall provided by the multi-scale cropping mechanism compensates for potential object boundary fragmentation. In contrast, multi-resolution methodsβthough preserving global structureβnegatively impact performance, likely due to the modelβs reliance on fixed-size input distributions during pretraining. Adjusting resolution without architectural adaptation may lead to degraded feature extraction in the frozen backbone of black-box MLLMs.
Backbone Variants for Region Extraction.
We also evaluate the effect of using different visual backbone models for region proposal and saliency detection. Specifically, we compare GroundingDINOΒ (liuGroundingDINOMarrying2023) and EVF-SAM as RoI extractors in the multi-scale visual emphasis pipeline. Both variants yield significant improvements over the baseline; however, GroundingDINO exhibits slightly higher accuracy across datasets. This result suggests that while both models effectively localize task-relevant regions, the localization precision and language grounding of GroundingDINO may offer a marginal advantage in aligning visual slices with text prompts.
Area Ratio Threshold
We vary the threshold in Eq.Β equationΒ 1 among and report results on the VSTAR validation set.
| Accuracy | Relative token usage | |
| 0.25 | 0.645 | 0.82 |
| 0.35 | 0.652 | 0.79 |
| 0.45 | 0.647 | 0.76 |
We observe that achieves the best trade-off between accuracy and token efficiency: smaller thresholds append the global view too often, increasing token usage with marginal accuracy gains, while larger thresholds skip the global view in cases that benefit from additional context.
Synergistic Effects.
Combining the multi-scale emphasis with the Patches variant of yields the strongest overall performance, validating the importance of both region-level granularity and architectural alignment. These findings highlight that visual prompting should be co-designed with content-aware pre-processing and prompt-space organization, especially under token-constrained inference regimes.
@@ 6 Conclusion
This paper presents , a novel visual prompting framework that addresses the challenge of preserving fine-grained visual detail under token constraints in black-box multimodal language models (MLLMs). By decomposing the prompting task into prompt-aware emphasis, spatially structured orchestration, and budget-aware input construction, enables effective visual grounding without access to model internals or architectural modifications.
Comprehensive evaluations across datasets such as Vstar, RealWorldQA, and TerraIncognita demonstrate that consistently improves task accuracy while reducing token consumption. For example, -Patches achieves a 26.9% gain over baseline accuracy on Vstar, while -Adaptive improves RealWorldQA accuracy by 12.1%. On TerraIncognita, achieves higher accuracy with 67% fewer tokens, highlighting its practical utility in resource-constrained settings.
While the present work focuses on token-level efficiency, an important emerging challenge is the communication cost associated with transferring high-resolution images from edge devices to cloud-based MLLMs. This concern is particularly relevant in scenarios such as wearable computing or mobile robotics. A promising direction for future research is to extend toward edge-aware visual prompting, enabling lightweight pre-processing directly on-device to reduce upstream bandwidth and latency while maintaining task performance.
In summary, contributes a modular and model-agnostic framework for enhancing visual processing in black-box MLLMs. Its integration of spatial structure, semantic relevance, and token budget constraints offers a principled foundation for both immediate deployment and future extensions in bandwidth-constrained or real-time vision-language applications.
From heuristics to explicit optimization.
The current Zoomer pipeline is heuristic in nature: it relies on prompt-aware detection, spatial-preserving composition, and simple token budgeting rules. A natural next step is to formalize adaptive visual token allocation as a constrained optimization problem, e.g., maximizing task accuracy or expected utility under a fixed visual-token budget and latency constraint, and to develop trainable surrogates or reinforcement learning policies that optimize this objective end-to-end.
Beyond external detectors.
Zoomer currently relies on open-vocabulary detectors (e.g., GroundingDINO, EVF-SAM) applied in a multi-scale fashion, which can become a bottleneck when detector performance degrades. Our experiments partially mitigate this via Zoomer-Adaptive and Zoomer-Global, which automatically include a global view when the cropped area ratio is small. Nonetheless, exploring detector-light or detector-free mechanismsβsuch as leveraging internal attention maps or saliency predictions from open-source MLLMsβis an exciting direction for future work.